OpenStudy (anonymous):

integrate 1/(sqrtx(4x+1))

ganeshie8 (ganeshie8):

try u^2 = 4x+1

OpenStudy (ahsome):

Is the equation: \[\frac{1}{\sqrt{x(4x+1)}}\]?

ganeshie8 (ganeshie8):

Oh looks i missed that x in front when I looked at it earlier :o

OpenStudy (ahsome):

I think it is, I am not sure

OpenStudy (anonymous):

yes that is the equation

OpenStudy (anonymous):

oh wait no only the first x is under the square root

OpenStudy (unklerhaukus):

\[\int\frac{\mathrm dx}{\sqrt x(4x+1)}\]

OpenStudy (anonymous):

yess unklerhaukus that is the correct equation

OpenStudy (ahsome):

Why is it \(dx\) and not 1?

OpenStudy (anonymous):

it's the same thing

OpenStudy (ahsome):

Oh, thanks for the clarrification.

OpenStudy (unklerhaukus):

try the substitution \(u=\sqrt x\) \(\mathrm d u = \,?\,\mathrm dx\)

OpenStudy (anonymous):

I don't think that works...du=1/2x^(-1/2) ??

OpenStudy (unklerhaukus):

\[\mathrm du =\frac12x^{-1/2} \mathrm dx\] that is correct , but if you write it another way \[\mathrm du = \frac{\mathrm dx}{2\sqrt x}\]

OpenStudy (unklerhaukus):

and you have \(\dfrac{\mathrm dx}{\sqrt x}\) in your integrand already, so it will cancel out nicely

OpenStudy (unklerhaukus):

What do you get for the integral in-terms of u and du

OpenStudy (anonymous):

im not sure what to sub for the (4x+1)

OpenStudy (unklerhaukus):

we choose \(u = \sqrt x\) so \(u^2 =x\)

OpenStudy (anonymous):

du/(4u^2+1)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

and then would i use partial fractions?

OpenStudy (unklerhaukus):

careful, dx/√x = 2du

OpenStudy (anonymous):

oh yeah i took the 2 outside the integrand

OpenStudy (unklerhaukus):

i think you might find a similar integral in a table

OpenStudy (unklerhaukus):

\[\boxed{\displaystyle\int\frac{\mathrm dx}{a^2+x^2} = \frac1a\arctan\tfrac xa+c}\]

OpenStudy (unklerhaukus):

you'll need you manipulate the integrand a bit until you have this form

OpenStudy (unklerhaukus):

@fgmel88 , how are you going with it?